the length of a rectangular mailing label is 5 cm less than twice the width. The perimeter is 20 cm. find the dimensions of the label
Answer by Jk22(389)

let l=length, w=width, p=perimeter (in cm)
data : p=20, l=2w-5
know : p=2l+2w=2l+l+5=3l+5 (warning : not to read as 21+1=22)
=> 20=3l+5 => l=5, w=(l+5)/2=5
hence it were a square of 5cm side.